Bella's brother is 9 years older than Bella. Their ages add up to 19. How old is Bella?
Answer by jim_thompson5910(35256)

x = bella's age, y = brother's age

y = x+9
x + y = 19

----------------------------------

x+y = 19

x+(x+9) = 19

x+x+9 = 19

2x+9 = 19

2x = 19-9

2x = 10

x = 10/2

x = 5
--------------------------
y = x+9

y = 5+9

y = 14

=======================================
So the answers are x = 5 and y = 14


So Bella is 5 years old and her brother is 14 years old

Check:

1) Make sure that her brother is indeed 9 years older. This is true since 14-5 = 9 (ie this shows that her brother is truly 9 years older)

2) Make sure their ages add to 19. So 5 + 14 = 19, which is also true

So both checks pass, which verify the answer.